OpenStudy (dls):

How do we solve 10^16-10^4?

OpenStudy (lgbasallote):

\[10^{16} - 14^{4}?\]

OpenStudy (dls):

its 10 sorry not 14

OpenStudy (dls):

and yes that only

OpenStudy (lgbasallote):

dont you just use a calculator? o.O

OpenStudy (dls):

how to do it without that?

OpenStudy (usukidoll):

that's one tough cookie . Need to use scientific notation to get those meanings or you can just do 1.6 x 10 ^ ???

OpenStudy (lgbasallote):

uhh you do it manually :/

OpenStudy (dls):

I just want a simpler answer not manual one!

OpenStudy (lgbasallote):

there is none -_- there are no algebra rules for \[x^m - x^n\]

OpenStudy (usukidoll):

that's what I'm trying to do . manually. since putting 10^16 is impossible in a calculator. A calculator can only handle 9 digits.

OpenStudy (lgbasallote):

unless they added one while i was sleeping in the cryogenic chamber...

OpenStudy (usukidoll):

10^4 is the easy part 10^16...hold up...

OpenStudy (dls):

we can't take 10 common?it wont help :o

OpenStudy (lgbasallote):

you can take 10^4 common

OpenStudy (usukidoll):

10 ^ 2 = 10 x 10 It's a real time consuming process to write 10 16 times

OpenStudy (lgbasallote):

\[10^4 (10^{12} - 1)\]

OpenStudy (lgbasallote):

but i dont think that will help...

OpenStudy (usukidoll):

scientific notation. I don't remember any factoring when dealing with those problems

OpenStudy (dls):

999999999999 x 10000 ! nvm

OpenStudy (dls):

9999999999990000 lol

OpenStudy (lgbasallote):

huh? how'd that become the answer o.O

OpenStudy (usukidoll):

ummm... 1 0 fifteen more zeros. omg where is my book lol. sorry it's been years since I last dealt with this.

OpenStudy (lgbasallote):

to wolfram i go..

OpenStudy (dls):

lol

OpenStudy (usukidoll):

10^4 = 10 x 10 x 10 x 10 10 ^ 16 = 10 x 10 x 10 x 10 x 10 x 10 x 10 x 10 x10 x 10 x 10 x 10 x10 x 10 x 10 x 10 <--- impossible to get the answer on a calculator.

OpenStudy (lgbasallote):

hmm yeah 9.99 x 10^15 lol

OpenStudy (lgbasallote):

but there's no simple way of doing it :/

OpenStudy (lgbasallote):

well 10^12 - 1 is 9.99 x 10^11

OpenStudy (usukidoll):

but it's 10 ^ 16 - 10 ^ 4 meaning 10^16 - 10,000

OpenStudy (dls):

thats what i did!

OpenStudy (lgbasallote):

9.99 x10^11 x 10^4 = 9.99 x10^15 i guess that is simple

OpenStudy (dls):

i guess that's the answer :P

OpenStudy (lgbasallote):

wolfram agrees

OpenStudy (usukidoll):

wolfram alpha I miss using that program

OpenStudy (lgbasallote):

i developed it. my middle name is wolfram.

OpenStudy (dls):

oh..!

OpenStudy (usukidoll):

I see... 

OpenStudy (usukidoll):

10000000000000000-10000

OpenStudy (dls):

bingo :P